OpenStudy (anonymous):

Solve for x 3 1 - - - =1 5x 2x

OpenStudy (phi):

Do you know how to make the denominator 10x? multiply the first fraction by 2/2 mult the 2nd fraction by 5/5 now you can subtract the tops

OpenStudy (phi):

\[ \frac{2}{2}\cdot\frac{3}{5x} -\frac{5}{5}\cdot\frac{1}{2x} \] \[ \frac{6}{10x} -\frac{5}{10x} \] \[ \frac{6-5}{10x} = \frac{1}{10x} \]